OpenStudy (anonymous):

In 3x + 2y = 23 x + y = 16 and 2x + 3y = 15 3x + 2y = 0 Do x = 6 Y = 10 and X = -6 Y=9

OpenStudy (anonymous):

u got the second solutions to the 2nd equation correct but not for the first one..:)

OpenStudy (anonymous):

for the first one x=-9 and y=25

OpenStudy (anonymous):

In the equations 3x + 2y = 23 x + y = 16 x = 6 y = 10 and in the equation 2x + 3y = 15 3x + 2y = 0 X = -6 Y=9

OpenStudy (anonymous):

MSFFL:see what i posted :)

OpenStudy (anonymous):

what about the second one?

OpenStudy (anonymous):

u r solutions for the 2nd one are correct! :)

OpenStudy (anonymous):

Phew. that means I passed my quiz in algebra!

OpenStudy (anonymous):

good luck..tc:)

OpenStudy (anonymous):

I only got two out of eight wrong. I both were careless mistakes though. Got be more careful.